A social psychologist wants to test if extroversion is significantly related to a person’s general health. He gets extroversion scores and a measure of general health from the participants’ medical providers. The data are listed below. Higher health extroversion scores mean greater extroversion, while high health scores = good health.

Complete the appropriate statistical test and hypothesis testing using the 4 step approach. State your conclusion in APA format. Alpha level = .05.

Participant

extrov X

health Y

X squared Y squared (X) (Y)
A

101

33 3333
B 93 40 3720
C 85 24 2040
D 111 46 5106
E 104 30 3120
F 89 28 2492
Total

Then,

Using the data from the prior correlation question (extroversion and health),

Find the regression line for the data.

a = ____       b = ______        equation:   _____________________
Answer #1

solution 8 = n=6 Ex-583 Ey= = 201 Exy = 19311 sst = 282 ssy = 310 sexy = {(x1 - x) lyi-y) = {xy-nxy 97.1667 ។ Ey 201 = 33.5 STest statistico rhoz - tnoz V1-82 1.8972 5-9992 = 0.948656-2 ſi-0-94962 0.3164 P-value=0.002 tcritical=tn-2 , 2/2 = t4, 0.025